# Coincidence detection techniques for direct measurement of many-body correlations in strongly correlated electron systems

### v1

@@ I Introduction
Strongly correlated electron systems represent a major focus of modern condensed matter physics due to their novel physical properties. Systems such as unconventional superconductors (Anderson, 1987; Lee et al., 2006; Chen et al., 2014; Stewart, 2011, 1984, 2001) and quantum magnets (Balents, 2010; Zhou et al., 2017; Chamorro et al., 2021) exhibit a range of unconventional and anomalous phenomena, including non-BCS superconductivity, pseudogap physics, strange metal behavior, electronic nematicity, quantum spin liquids, and quantum phase transitions. These emergent properties are governed by intrinsic many-body correlations that can not be adequately described within well established theoretical frameworks like the perturbative Fermi liquid theory or the Ginzburg–Landau-Wilson theory of phase transitions. The development of new theoretical and experimental methods to study these correlation-driven phenomena remains a significant challenge. In this perspective article, we focus on recent progress and ongoing efforts to develop new experimental techniques within this research field.
@@
As outlined in the above brief review, a wide array of experimental techniques has been employed to investigate strongly correlated electron systems. Despite the substantial body of experimental data thus accumulated, research in this field appears to be approaching a bottleneck. This impasse may be attributed to two principal factors. First, the essential intrinsic signatures have yet to be systematically distilled from the vast and complex experimental data, hindering the development of sound phenomenological theories. Second, the existing experimental approaches may be fundamentally limited in their ability to probe the intrinsic many-body correlations and their physical manifestations. Unraveling the foundational correlation-driven physics will likely require the development of new experimental theories and advanced technologies.
@@
The intrinsic $N$-body correlations in strongly correlated electron systems can be accessed via $N$-body response measurements. For example, the intrinsic two-body correlations can be directly probed via two-body response functions which are dominated by the second-order perturbation of the interaction between the target material and an external probe field. Guided by this principle, several coincidence detection techniques have recently been proposed for directly measuring two-body correlations in strongly correlated electron systems. Notable examples include coincidence angle-resolved photoemission spectroscopy (cARPES), which is designed to access two-body correlations in the particle-particle channel, thereby offering a new pathway to elucidate the mechanism of unconventional superconductivity (Su and Zhang, 2020; Devereaux et al., 2023; Cao and Su, 2025). Meanwhile, coincidence inelastic neutron scattering (cINS) has been proposed to probe two-spin correlations, opening avenues for studying exotic quantum spin liquids (Su et al., 2021; Cao and Su, 2025). More recently, a coincidence double-tip scanning tunneling spectroscopy (STS) method has also been introduced, enabling spatially resolved detection of dynamical two-body correlations (Su et al., 2025). Other proposed coincidence detection techniques include coincidence angle-resolved photoemission and inverse-photoemission spectroscopy (cARP/IPES) in the particle-hole channel for studying itinerant magnetism and electronic nematicity (Su and Zhang, 2020), coincidence angle-resolved inverse-photoemission spectroscopy (cARIPES) in the particle-particle channel for correlated electrons above the Fermi energy (Su and Zhang, 2020), and double photoemission (DPE) for exploring center-of-mass physics of two-body correlations (Berakdar, 1998; Su et al., 2024). The fundamental principles underlying these coincidence detection techniques can be summarized as follows (Su et al., 2024): (1) Intrinsic two-body correlations are directly captured through two-body coincidence detection, and (2) the coincidence detection probability is governed by the second-order perturbation of the interaction between the target material and the external probe field.
@@ III Coincidence Detection Techniques
Coincidence detection of two photoelectric processes, in which two incident photons emit two photoelectrons from the target material, captures two-body correlations of the correlated electrons in the particle-particle channel. — cARPES
In cARPES coincidence detection (schematically illustrated in Fig. 1 (a) of Reference (Cao and Su, 2025)), two incident photons induce two photoelectric processes. Two single-photoelectron detectors separately detect the emitted photoelectrons. A coincidence detector records the simultaneous occurrence events of two photoelectric processes. The corresponding coincidence detection probability carries information on two-body correlations of the target electron system.
@@
The coincidence detection probability for the simultaneous occurrence of two photoelectric processes is given by (Su and Zhang, 2020)
$\Gamma^{(2)}=\frac{1}{Z}\sum_{\alpha\beta}e^{-\beta E_{\alpha}}\left\lvert\langle\Phi_{\beta}\rvert S^{(2)}(t_{f},t_{i})\rvert\Phi_{\alpha}\rangle\right\rvert^{2},$ (14)
with initial states $\lvert\Phi_{\alpha}\rangle=\lvert\Psi_{\alpha}\rangle\otimes\lvert 1_{\mathbf{q}_{1}\lambda_{1}}1_{\mathbf{q}_{2}\lambda_{2}}\rangle_{p}\otimes\lvert 0\rangle_{d}$ and final states $\lvert\Phi_{\beta}\rangle=\lvert\Psi_{\beta}\rangle\otimes\lvert 0\rangle_{p}\otimes\lvert 1_{\mathbf{k}_{1}+\mathbf{q}_{1}\sigma_{1}}1_{\mathbf{k}_{2}+\mathbf{q}_{2}\sigma_{2}}\rangle_{d}$. Here $\lvert\Psi_{\alpha}\rangle$ and $\lvert\Psi_{\beta}\rangle$ are eigenstates of the target electron system with energies $E_{\alpha}$ and $E_{\beta}$, respectively. The relevant second-order $S$-matrix term reads
$S^{(2)}(t_{f},t_{i})=\left(-\frac{i}{\hbar}\right)^{2}\iint^{t_{f}}_{t_{i}}dt_{2}dt_{1}T_{t}\left[V_{2I}(t_{2})V_{1I}(t_{1})\right].$ (15)
@@
In the limit $t_{i}\to$-$\infty$ and $t_{f}\to+\infty$, the coincidence detection probability becomes (Su and Zhang, 2020)
@@
Here we have identified $\phi^{(2)}_{\alpha\beta}\left(\mathbf{k}_{1}\sigma_{1},\mathbf{k}_{2}\sigma_{2};t_{c},t_{r}\right)=\phi^{(2)}_{\alpha\beta}\left(\mathbf{k}_{1}\sigma_{1}t_{1};\mathbf{k}_{2}\sigma_{2}t_{2}\right)$, where $t_{c}=\frac{1}{2}(t_{1}+t_{2})$ and $t_{r}=t_{2}-t_{1}$ denote the center-of-mass and relative times, respectively. The corresponding Fourier transformations are defined by
$\phi^{(2)}_{\alpha\beta}\left(\mathbf{k}_{1}\sigma_{1},\mathbf{k}_{2}\sigma_{2};t_{c},t_{r}\right)$
$=$ $\frac{1}{\left(2\pi\right)^{2}}\iint_{-\infty}^{+\infty}d\Omega d\omega\,\phi^{(2)}_{\alpha\beta}\left(\mathbf{k}_{1}\sigma_{1},\mathbf{k}_{2}\sigma_{2};\Omega,\omega\right)e^{-i\Omega t_{c}-i\omega t_{r}},$
$\phi^{(2)}_{\alpha\beta}\left(\mathbf{k}_{1}\sigma_{1},\mathbf{k}_{2}\sigma_{2};\Omega,\omega\right)$
$=$ $\iint_{-\infty}^{+\infty}dt_{c}dt_{r}\,\phi^{(2)}_{\alpha\beta}\left(\mathbf{k}_{1}\sigma_{1},\mathbf{k}_{2}\sigma_{2};t_{c},t_{r}\right)e^{i\Omega t_{c}+i\omega t_{r}}.$
The center-of-mass frequency $\Omega_{c}$ and the inner-pair relative frequency $\omega_{r}$ in $\Gamma^{(2)}$ are defined as
@@
with $E_{1}^{(2)}$ and $E_{2}^{(2)}$ being the transferred energies in the two photoelectric processes
@@
Here $\varepsilon^{(d)}$ are the photoelectron energies, $\hbar\omega$ are the photon energies, and $\Phi$ is the target-material work function.
@@
In the initial experimental design for cARPES measurement (Su and Zhang, 2020; Cao and Su, 2025), a coincidence detector is employed to instantaneously record the simultaneous occurrence of two photoelectric processes. This configuration is termed instantaneous cARPES. More recently, an alternative implementation was proposed (schematically shown in Fig. 1 (b) of Reference (Cao and Su, 2025)), termed post-experiment cARPES (Cao and Su, 2025). To capture the dynamical two-body correlations in the target electron system, the time interval between the initiations of two photoelectric processes must be shorter than the characteristic time scale $t_{c}$ of the physics under investigation. This can be achieved using a pulse photon source, where both incident photons originate from the same pulse with a time width $\Delta t_{p}\leq t_{c}$. Furthermore, to clearly distinguish individual coincidence detection events, the time interval $\Delta t_{d}$ between consecutive photon pulses must satisfy $\Delta t_{d}\gg\Delta t_{p},t_{c}$. A further refinement in the post-experiment cARPES is the use of two counting recorders $R_{1}$ and $R_{2}$, which record photoelectrons arriving at two detectors $D_{1}$ and $D_{2}$, respectively. For the $n$-th photon pulse, the recorded photoelectron numbers $a_{1}(n)$ and $a_{2}(n)$ in $R_{1}$ and $R_{2}$ take values of 0 or 1. The coincidence counting for the $n$-th pulse is then computed as $a_{1}(n)\times a_{2}(n)$ during post-processing. This approach, without the coincidence detector used in the instantaneous cARPES, enables easier and more efficient identification of two-photoelectric-process coincidence events. As a result, the post-experiment cARPES technique offers a highly efficient and powerful means for measuring two-body correlations in the target electron system (Cao and Su, 2025).
It is worth noting that the cARPES technique shares conceptual similarities with time-resolved ARPES measurements of two-electron emission events, which was proposed to investigate electronic noise correlations in materials (Stahl and Eckstein, 2019). The related theoretical framework has also been developed to probe two-electron interactions via coincidence detection of two photoelectric processes (Kemper et al., 2025).
@@
Following a procedure analogous to that of cARPES, we can introduce the second-order term of the time-evolution $S$-matrix corresponding to the two electron-neutron interaction terms, enabling the calculation of the coincidence detection probability $\Gamma^{(2)}$ for the simultaneous occurrence of two neutron-scattering processes. Let us define a two-spin Bethe-Salpeter wave function as (Su et al., 2021; Cao and Su, 2025)
@@
where $i$ and $j$ label the Cartesian components of the spin operator. It can be shown that the coincidence detection probability provides direct information on $\phi^{(ij)}_{\alpha\beta}(\mathbf{q}_{1},\mathbf{q}_{2};\Omega_{c},\omega_{r})$, which is obtained via Fourier transformation of $\phi^{(ij)}_{\alpha\beta}(\mathbf{q}_{1},\mathbf{q}_{2};t_{c},t_{r})=\phi^{(ij)}_{\alpha\beta}(\mathbf{q}_{1}t_{1},\mathbf{q}_{2}t_{2})$.
@@
Here, the creation and annihilation operators are defined as in Eq. (13).
@@ III.4 Coincidence double-tip STS
A coincidence detection technique for directly probing spatially resolved dynamical two-body correlations has recently been proposed, based on the double-tip scanning tunneling microscope (STM) and termed coincidence double-tip STS (Su et al., 2025). A schematic illustration of its setup is provided in Fig. 1 of Reference (Su et al., 2025).
@@
Experimental implementation of the double-tip STS has advanced considerably over the past two decades (Shiraki et al., 2001; Nakayama et al., 2012; Voigtländer et al., 2018; Leeuwenhoek et al., 2020a, b). However, achieving precise control over tip positionings at extremely short length scales remains a major challenge. Modern double-tip STM setups have successfully realized tip separations on the order of $\sim 30~\text{nm}$ (Nakayama et al., 2012; Leeuwenhoek et al., 2020b). Nevertheless, such resolutions are still insufficient for studying most strongly correlated electron systems, such as various classes of unconventional superconductors (Lee et al., 2006; Chen et al., 2014; Stewart, 2011), where dominant strong correlations are highly localized, with typical correlation lengths comparable to interatomic distances (a few Å). The recent discovery of unconventional superconductivity and Mott physics in magic-angle twisted bilayer graphene (Cao et al., 2018) offers a promising alternative platform for coincidence double-tip STS. In this system, exotic correlated phenomena are inherently connected to the moiré superlattice, which exhibits a characteristic periodicity of $\sim 13~\text{nm}$, a mesoscopic length scale within the resolution of existing double-tip STM setups. Consequently, moiré quantum materials represent an ideal platform for implementing coincidence double-tip STS to probe spatially resolved dynamical two-body correlations in strongly correlated electron systems.
@@ IV Discussion and Future Prospects
Looking ahead, the development of novel coincidence detection techniques specifically tailored for strongly correlated electron systems holds great promise. Moreover, given the strong entanglement among spin, charge, and orbital degrees of freedom in such materials, approaches capable of directly probing correlations between these different degrees of freedom could open entirely new research directions. In particular, the simultaneous detection of two physical processes involving distinct degrees of freedom would allow direct measurement of their mutual correlations and entanglement. Such methods hold the potential to greatly deepen our understanding of strongly correlated quantum matters, especially in shedding light on the complex intertwined quantum phases arising from such strong entanglement.

### v2

@@ I Introduction
Strongly correlated electron systems represent a major focus of modern condensed matter physics due to their novel physical properties. Systems such as unconventional superconductors (Anderson, 1987; Lee et al., 2006; Chen et al., 2014; Stewart, 2011, 1984, 2001) and quantum magnets (Balents, 2010; Zhou et al., 2017; Chamorro et al., 2021) exhibit a range of unconventional and anomalous phenomena, including non-BCS superconductivity, pseudogap physics, strange metal behavior, electronic nematicity, quantum spin liquids, and quantum phase transitions. These emergent properties are governed by intrinsic many-body correlations that can not be adequately described within well established theoretical frameworks like the perturbative Fermi liquid theory or the Ginzburg-Landau-Wilson theory of phase transitions. The development of new theoretical and experimental methods to study these correlation-driven phenomena remains a significant challenge. In this perspective article, we focus on recent progress and ongoing efforts to develop new experimental techniques within this research field.
@@
As outlined in the above brief review, a wide array of experimental techniques has been employed to investigate strongly correlated electron systems. Despite the substantial body of experimental data thus accumulated, research in this field appears to be approaching a bottleneck. This impasse may be attributed to two principal factors. First, the essential intrinsic signatures have yet to be systematically distilled from the vast and complex experimental data, hindering the development of sound phenomenological theories. Second, the existing experimental approaches may be fundamentally limited in their ability to probe the intrinsic many-body correlations and their physical manifestations. Breaking through this impasse and unraveling the foundational correlation-driven physics will likely require the development of new experimental theories and advanced technologies.
Figure 1: Schematic illustration of coincidence detection. (a) Fundamental idea: Coincidence detection of two single-particle perturbation processes reveals dynamical two-body correlations. (b) By coincidence detection of two photoelectric processes, coincidence angle-resolved photoemission spectroscopy (cARPES) captures the dynamical two-body correlations in the particle-particle channel. (c) By coincidence detection of two neutron-scattering processes, coincidence inelastic neutron scattering (cINS) uncovers dynamical two-spin correlations. (d) By coincidence detection of two quantum tunneling currents, coincidence double-tip scanning tunneling spectroscopy (STS) probes the spatially resolved dynamical two-body correlations. (e) Coincidence detection technique with post-experiment coincidence counting method. Other coincidence detection techniques, such as coincidence angle-resolved photoemission and inverse-photoemission spectroscopy (cARP/IPES), coincidence angle-resolved inverse-photoemission spectroscopy (cARIPES), and double photoemission (DPE), share a similar configuration to (a), but with different incident/scattered particle sources or detectors, for the dynamical two-body correlations in the different particle-particle or particle-hole channel.
@@
The intrinsic $N$-body correlations in strongly correlated electron systems can be accessed via $N$-body response measurements. For example, the intrinsic two-body correlations can be directly probed via two-body response functions which are dominated by the second-order perturbation of the interaction between the target material and an external probe field. Guided by this principle, several coincidence detection techniques have recently been proposed for directly measuring two-body correlations in strongly correlated electron systems, as schematically illustrated in Fig. 1. Notable examples include coincidence angle-resolved photoemission spectroscopy (cARPES), which is designed to access two-body correlations in the particle-particle channel, thereby offering a new pathway to elucidate the mechanism of unconventional superconductivity (Su and Zhang, 2020; Devereaux et al., 2023; Cao and Su, 2025). Meanwhile, coincidence inelastic neutron scattering (cINS) has been proposed to probe two-spin correlations, opening avenues for studying exotic quantum spin liquids (Su et al., 2021; Cao and Su, 2025). More recently, a coincidence double-tip scanning tunneling spectroscopy (STS) method has also been introduced, enabling spatially resolved detection of dynamical two-body correlations (Su et al., 2025). Other proposed coincidence detection techniques include coincidence angle-resolved photoemission and inverse-photoemission spectroscopy (cARP/IPES) in the particle-hole channel for studying itinerant magnetism and electronic nematicity (Su and Zhang, 2020), coincidence angle-resolved inverse-photoemission spectroscopy (cARIPES) in the particle-particle channel for correlated electrons above the Fermi energy (Su and Zhang, 2020), and double photoemission (DPE) for exploring center-of-mass physics of two-body correlations (Berakdar, 1998; Su et al., 2024). The fundamental principles underlying these coincidence detection techniques can be summarized as follows (Su et al., 2024): (1) Intrinsic two-body correlations are directly captured through two-body coincidence detection, and (2) the coincidence detection probability is governed by the second-order perturbation of the interaction between the target material and the external probe field.
@@ III Coincidence Detection Techniques
Coincidence detection of two photoelectric processes, in which two incident photons emit two photoelectrons from the target material, captures two-body correlations of the correlated electrons in the particle-particle channel.cARPES
In cARPES coincidence detection, as schematically illustrated in Fig. 1 (b), two incident photons induce two photoelectric processes. A coincidence detector records the simultaneous occurrence events of two photoelectric processes. The corresponding coincidence detection probability carries information on two-body correlations of the target electron system.
@@
where $\lvert\Psi_{\alpha}\rangle$ and $\lvert\Psi_{\beta}\rangle$ are eigenstates of the target electron system with energies $E_{\alpha}$ and $E_{\beta}$, respectively. In the limit $t_{i}\to$-$\infty$ and $t_{f}\to+\infty$, the coincidence detection probability for the simultaneous occurrence of two photoelectric processes follows (Su and Zhang, 2020)
@@
Here $\phi^{(2)}_{\alpha\beta}\left(\mathbf{k}_{1}\sigma_{1},\mathbf{k}_{2}\sigma_{2};\Omega_{c},\omega_{r}\right)$ is the frequency-space Bethe-Salpeter wave function. It is the Fourier transformation of $\phi^{(2)}_{\alpha\beta}\left(\mathbf{k}_{1}\sigma_{1},\mathbf{k}_{2}\sigma_{2};t_{c},t_{r}\right)=\phi^{(2)}_{\alpha\beta}\left(\mathbf{k}_{1}\sigma_{1}t_{1};\mathbf{k}_{2}\sigma_{2}t_{2}\right)$, where $t_{c}=\frac{1}{2}(t_{1}+t_{2})$ and $t_{r}=t_{2}-t_{1}$ denote the center-of-mass and relative times, respectively. The corresponding center-of-mass frequency $\Omega_{c}$ and the inner-pair relative frequency $\omega_{r}$ in $\Gamma^{(2)}$ are defined as
@@
where $E_{1}^{(2)}$ and $E_{2}^{(2)}$ are the transferred energies in the two photoelectric processes
@@
Here $\varepsilon^{(d)}_{\mathbf{k}+\mathbf{q}\sigma}$ are the photoelectron energies with momentum $\mathbf{k}+\mathbf{q}$ and spin $\sigma$, $\hbar\omega_{\mathbf{q}}$ are the photon energies with momentum $\mathbf{q}$, and $\Phi$ is the target-material work function.
@@
In the initial experimental design for cARPES measurement (Su and Zhang, 2020; Cao and Su, 2025), a coincidence detector is employed to instantaneously record the simultaneous occurrence of two photoelectric processes. This configuration is termed instantaneous cARPES. More recently, an alternative implementation was proposed (schematically shown in Fig. 1 (e)), termed post-experiment cARPES (Cao and Su, 2025). To capture the dynamical two-body correlations in the target electron system, the time interval between the initiations of two photoelectric processes must be shorter than the characteristic time scale $t_{c}$ of the physics under investigation. This can be achieved using a pulse photon source, where both incident photons originate from the same pulse with a time width $\Delta t_{p}\leq t_{c}$. Furthermore, to clearly distinguish individual coincidence detection events, the time interval $\Delta t_{d}$ between consecutive photon pulses must satisfy $\Delta t_{d}\gg\Delta t_{p},t_{c}$. A further refinement in the post-experiment cARPES is the use of two counting recorders $R_{1}$ and $R_{2}$, which record photoelectrons arriving at two detectors $D_{1}$ and $D_{2}$, respectively. For the $n$-th photon pulse, the recorded photoelectron numbers $a_{1}(n)$ and $a_{2}(n)$ in $R_{1}$ and $R_{2}$ take values of 0 or 1. The coincidence counting for the $n$-th pulse is then computed as $a_{1}(n)\times a_{2}(n)$ during post-processing. This approach, without the coincidence detector used in the instantaneous cARPES, enables easier and more efficient identification of two-photoelectric-process coincidence events. As a result, the post-experiment cARPES technique offers a highly efficient and powerful mean for measuring two-body correlations in the target electron system (Cao and Su, 2025).
It is worth noting that the cARPES technique shares conceptual similarity with time-resolved correlation ARPES measurements of two-electron emission events (Stahl and Eckstein, 2019). Both techniques have the similar fundamental formulations based on perturbation S-matrix theory (Su and Zhang, 2020; Devereaux et al., 2023; Cao and Su, 2025). While cARPES measurements probe dynamical two-body correlations in the equilibrium states, time-resolved correlation ARPES measurements investigate these correlations in time-evolution nonequilibrium states. By incorporating the temporal envelope of ultrashort photon pulses, time-resolved correlation ARPES measurements enable time-resolved noise two-body correlations in nonequilibrium states. This contrasts with cARPES measurements, which provide dynamical two-body correlations in equilibrium states within frequency space, resolving both center-of-mass and relative frequency dependencies. More recently, a variant termed correlation ARPES technique has been developed from its time-resolved predecessor (Kemper et al., 2025). While it has the same theoretical formulation as the cARPES technique (Su and Zhang, 2020; Devereaux et al., 2023; Cao and Su, 2025), this approach emphasizes the extraction of information on electron-electron interactions in weakly correlated electron systems through the coincidence detection of two photoelectric processes.
@@
Let us define a two-spin Bethe-Salpeter wave function as (Su et al., 2021; Cao and Su, 2025)
@@
where $i$ and $j$ label the Cartesian components of the spin operator. It can be shown that the coincidence detection probability $\Gamma^{(2)}$ for the simultaneous occurrence of two neutron-scattering processes provides direct information on $\phi^{(ij)}_{\alpha\beta}(\mathbf{q}_{1},\mathbf{q}_{2};\Omega_{c},\omega_{r})$, which is obtained via Fourier transformation of $\phi^{(ij)}_{\alpha\beta}(\mathbf{q}_{1},\mathbf{q}_{2};t_{c},t_{r})=\phi^{(ij)}_{\alpha\beta}(\mathbf{q}_{1}t_{1},\mathbf{q}_{2}t_{2})$.
@@
Here, the creation and annihilation operators and the coupling constants $g_{1}$ and $g_{2}$ are defined as in Eq. (13).
@@ III.4 Coincidence double-tip STS
A coincidence detection technique for directly probing spatially resolved dynamical two-body correlations has recently been proposed, based on the double-tip scanning tunneling microscope (STM) and termed coincidence double-tip STS (Su et al., 2025). A schematic illustration of its setup is provided in Fig. 1 (d).
@@
Experimental implementation of the double-tip STS has advanced considerably over the past two decades (Shiraki et al., 2001; Nakayama et al., 2012; Voigtländer et al., 2018; Leeuwenhoek et al., 2020a, b). However, achieving precise control over tip positionings at extremely short length scales remains a major challenge. Modern double-tip STM setups have successfully realized tip separations on the order of $30~\text{nm}$ (Nakayama et al., 2012; Leeuwenhoek et al., 2020b). Nevertheless, such resolutions are still insufficient for studying most strongly correlated electron systems, such as various classes of unconventional superconductors (Lee et al., 2006; Chen et al., 2014; Stewart, 2011), where dominant strong correlations are highly localized, with typical correlation lengths comparable to interatomic distances (a few angstroms). The recent discovery of unconventional superconductivity and Mott physics in magic-angle twisted bilayer graphene (Cao et al., 2018) offers a promising alternative platform for coincidence double-tip STS. In this system, exotic correlated phenomena are inherently connected to the moiré superlattice, which exhibits a characteristic periodicity of approximately $13~\text{nm}$, a mesoscopic length scale within the resolution of existing double-tip STM setups. Consequently, moiré quantum materials represent an ideal platform for implementing coincidence double-tip STS to probe spatially resolved dynamical two-body correlations in strongly correlated electron systems.
@@ IV Discussion and Future Prospects
Looking ahead, the development of novel coincidence detection techniques specifically tailored for strongly correlated electron systems holds great promise. Moreover, given the strong entanglement among spin, charge, and orbital degrees of freedom in such materials, approaches capable of directly probing correlations between these different degrees of freedom could open entirely new research directions. In particular, the coincidence detection of two physical processes involving distinct degrees of freedom would allow direct measurement of their mutual correlations and entanglement. Such methods hold the potential to greatly deepen our understanding of strongly correlated quantum matters, especially in shedding light on the complex intertwined quantum phases arising from such strong entanglement.